A deli owner is introducing two new menu items, A and B. The owner thinks 30% of customers will order item A. She thinks customers who order item A will also order item B 70% of the time and that customers who do not order item A are likely to order item B 40% of the time. What is the probability that a customer will order at least one new item?
Answer #1

P(A) = 0.3

P(B|A) = 0.7

P(B|A') = 0.4

P(A and B) = 0.3*0.7 = 0.21

P(B|A') = P(B) - P(A and B)

P(B) = 0.4 - 0.21 = 0.19

P(A or B) = P(A) + P(B) - P(A and B)

= 0.3 + 0.19 - 0.21

= 0.28

Ans: 0.28